OpenStudy (anonymous):

Simplify. -19+13+(-22)

OpenStudy (anonymous):

Remember : +(-) = - So, -19 + 13 - 22 Can you solve it now or not??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What is the problem you are facing??

OpenStudy (anonymous):

Do you know 19 + 22 = What??

OpenStudy (anonymous):

-19+-22= -41 -14+13= -28 -28/2= -14 answer is -14

OpenStudy (anonymous):

meant -41+13= -28 -28/2= -14

OpenStudy (anonymous):

@dezz22 Firstly don't solve the problem from your own and let others do it on their own.. And secondly, you are doing it wrong...

OpenStudy (anonymous):

@avaldez1 I am asking you something why are you looking for answer only?